# LongCat-Flash-Thinking-2601 Technical Report

### v1

@@ 1 Introduction
In this work, we introduce LongCat-Flash-Thinking-2601, a powerful and efficient Mixture-of-Experts (MoE) reasoning model with 560B total parameters and 27B activated parameters on average per token, featuring strong agentic reasoning capability. The pre-training of LongCat-Flash-Thinking-2601 largely follows the recipe of LongCat-Flash-Chat (Team et al., 2025a), retaining the original data distribution to preserve competitive general reasoning performance. Building on this foundation, we further extend the model toward large-scale agentic reasoning through a carefully designed mid-training stage. Compared to traditional reasoning, agentic behaviors typically involve long-horizon trajectories with proactive tool invocations. However, such interaction patterns are extremely scarce in real-world corpora, where most data consist primarily of natural language. As a result, a raw model is largely unfamiliar with agentic interaction dynamics, leading to inefficient exploration during the reinforcement learning stage. To address this challenge, we expose the model to moderate-scale synthesized structured agentic trajectories during mid-training, providing a strong initialization for agentic reasoning behaviors.
@@ 2 Pre-Training
Our model builds upon the pretraining recipe of LongCat-Flash-Chat (Team et al., 2025a), inheriting its data distribution to preserve strong general-purpose language and reasoning capabilities. Beyond traditional reasoning, agentic behaviors typically involve long-horizon trajectories with proactive tool invocations. This brings two additional challenges: (i) substantially increased demands on long-context modeling efficiency, and (ii) the scarcity of large-scale agentic trajectories in the real world. To address the long-context requirements of agentic reasoning, we adopt a staged mid-training procedure with progressively increasing context lengths, allocating 500B tokens to the 32K/128K stages and an additional 40B tokens to the 256K stage. Since large-scale reinforcement learning alone is often inefficient and unstable without a model that has already been primed with basic agentic behaviors, we choose to expose the model to moderate-scale agentic data at this stage. To mitigate the scarcity of real-world agentic trajectories, we construct a hybrid data synthesis pipeline to collect and construct agentic training data. Furthermore, we predict optimal hyperparameters in the mid-training stage based on any pre-trained checkpoint, specifically designed to minimize the computational cost of finding the best configuration (see Appendix A).
@@ 3 Scaling Reinforcement Learning
Post-training with reinforcement learning has become a primary approach to elicit stronger reasoning capabilities. Following LongCat-Flash-Thinking (Team et al., 2025c), we adopt a unified multi-domain post-training pipeline, in which domain-specialized expert models are first trained under a shared framework and then consolidated into a single general model through both model-level and data-level merging.
@@ 3.1.2 Initial Policy
For agentic search capability, constructing high-quality trajectories requires explicitly modeling multi-step evidence gathering and complete condition verification, while avoiding spurious shortcuts based on partial information. Such structured reasoning traces are rarely available in real-world search logs. We therefore construct synthetic reasoning trajectories that prioritize correctness, reasoning completeness, and robustness against shortcut behaviors. We therefore construct synthetic reasoning trajectories that prioritize correctness, reasoning completeness, and robustness against spurious shortcuts. We apply strict filtering to ensure trajectory quality, removing trivial cases and enforcing a consistent reasoning and tool-use format. To avoid shortcut learning such as lucky guesses based on partial evidence, we require trajectories contains explicitly verification of all conditions specified in the query. We further retain trajectories that involve long and iterative exploration by compressing earlier steps, allowing long-horizon reasoning to be preserved without length constraints. Finally, we scale data collection and increase behavioral diversity by reusing rollout trajectories from subsequent reinforcement learning stages.
@@ 3.2 Scalable Asynchronous Agentic RL Framework
Compared with single-turn scenarios such as standard reasoning tasks, agentic training involves multi-turn interactions with variable environments (ENV) or tools, which pose new challenges to RL Infrastructure. During a typical multi-turn rollout stage, it repeatedly interleaves the LLM generation, environment execution, and reward assessment to construct the final trajectory for RL training. In this setting, trajectories are not only long-tailed skewed but also involve unpredictable and latency-skewed environment interactions, which lead to device underutilization under batch setting (Fu et al., 2025; Zhang et al., 2025a; Lu et al., 2025b). Additionally, our production cluster consists of mid-range accelerators, especially with only around 60GB of available device memory. The hardware constraints, together with its software ecosystem, pose significant challenges to achieving stable and scalable agentic RL training. To address these problems, we extend our multi-version asynchronous training system, DORA (Dynamic ORchestration for Asynchronous rollout) (Team et al., 2025b), to fully support large-scale RL training in multi-turn agentic scenarios.
@@
To minimize device idleness in the agentic setting, we introduce a fully streaming asynchronous pipeline both within the rollout process and between rollout and training, based on streaming RPC (Team et al., 2025b). Within the rollout loop in the RolloutManager, we remove batch barriers, enabling LLM generation, environment execution, and reward computation to be executed on remote workers at the granularity of individual samples. This prevents accelerators from idling while waiting for batched ENV calls to complete. To further address the long-tailed generation problem with training stability, our DORA system enables multi-version asynchronous training, where trajectories generated by different model versions are immediately enqueued upon completion. Within a step, our multi-version generation instances continue rollout using multiple previous model versions, while the Trainer can initiate training as soon as its conditions are met, or elastically scale up additional generation instances for free extra throughput when training devices are idle.
@@ 7 Conclusion
Dengchang Zhao Jiarui Zhao Shijun Dai Yang Bai Yuxin Chen
Di Xiu Jingang Wang Shuaikang Liu Yang Liu Yuxin Liu
Dishan Liu Jinluan Yang Shuang Zhou Yang Yang Zeyang Yu
Dongyu Ru Jinrui Ding Shuo Wang Yang Zheng Zhao Yang
Dunwei Tu Jinwei Xiao Songyuan Zhao Yanyu Chen Zhengsheng Huang
Fan Wu Jiyuan He Tao Liang Yaoming Wang Zhengyu Chen
Fengcheng Yuan Juncan Xu Tianhao Hu Yaoming Zhu Zhijian Liu
Fengcun Li Kefeng Zhang Tianze Chen Yaorui Shi Zhikang Xia
Gang Xu Keheng Wang Wei Liu Yaqi Huo Zhimin Lin
Guanyu Wu Li Wei Wei Shi Yerui Sun Zhiyuan Yao
Guoyuan Lin Lianhui Ma Wei Wang Yi Zhang Zhuofan Chen
Haibin Wang Lin Qiu Weifeng Tang Yifan Lu Zhuowen Han
Hansi Yang Lingbing Kong Wenjie Shi Yifan Zhao Zijian Zhang
Hao Yang Lingchuan Liu Wenlong Zhu Yihao Chen Ziran Li
Haonan Yan Linsen Guo Wentao Chen Yi-Kai Zhang Ziwen Wang
Haoxiang Ma Mengshen Zhu Wentao Shi* Yitao Zhai Ziyuan Zhuang
Haoxing Wen Mengxia Shen Xi Su Yongjing Yin LongCat-Flash

### v2

@@ 1 Introduction
In this work, we introduce LongCat-Flash-Thinking-2601, a powerful and efficient Mixture-of-Experts (MoE) reasoning model with 560B total parameters and 27B activated parameters on average per token, featuring strong agentic reasoning capability. The pre-training of LongCat-Flash-Thinking-2601 largely follows the recipe of LongCat-Flash-Chat (Meituan et al., 2025), retaining the original data distribution to preserve competitive general reasoning performance. Building on this foundation, we further extend the model toward large-scale agentic reasoning through a carefully designed mid-training stage. Compared to traditional reasoning, agentic behaviors typically involve long-horizon trajectories with proactive tool invocations. However, such interaction patterns are extremely scarce in real-world corpora, where most data consist primarily of natural language. As a result, a raw model is largely unfamiliar with agentic interaction dynamics, leading to inefficient exploration during the reinforcement learning stage. To address this challenge, we expose the model to moderate-scale synthesized structured agentic trajectories during mid-training, providing a strong initialization for agentic reasoning behaviors.
@@ 2 Pre-Training
Our model builds upon the pretraining recipe of LongCat-Flash-Chat (Meituan et al., 2025), inheriting its data distribution to preserve strong general-purpose language and reasoning capabilities. Beyond traditional reasoning, agentic behaviors typically involve long-horizon trajectories with proactive tool invocations. This brings two additional challenges: (i) substantially increased demands on long-context modeling efficiency, and (ii) the scarcity of large-scale agentic trajectories in the real world. To address the long-context requirements of agentic reasoning, we adopt a staged mid-training procedure with progressively increasing context lengths, allocating 500B tokens to the 32K/128K stages and an additional 40B tokens to the 256K stage. Since large-scale reinforcement learning alone is often inefficient and unstable without a model that has already been primed with basic agentic behaviors, we choose to expose the model to moderate-scale agentic data at this stage. To mitigate the scarcity of real-world agentic trajectories, we construct a hybrid data synthesis pipeline to collect and construct agentic training data. Furthermore, we predict optimal hyperparameters in the mid-training stage based on any pre-trained checkpoint, specifically designed to minimize the computational cost of finding the best configuration (see Appendix A).
@@ 3 Scaling Reinforcement Learning
Post-training with reinforcement learning has become a primary approach to elicit stronger reasoning capabilities. Following LongCat-Flash-Thinking (Meituan. et al., 2025), we adopt a unified multi-domain post-training pipeline, in which domain-specialized expert models are first trained under a shared framework and then consolidated into a single general model through both model-level and data-level merging.
@@ 3.1.2 Initial Policy
For agentic search capability, constructing high-quality trajectories requires explicitly modeling multi-step evidence gathering and complete condition verification, while avoiding spurious shortcuts based on partial information. Such structured reasoning traces are rarely available in real-world search logs. We therefore construct synthetic reasoning trajectories that prioritize correctness, reasoning completeness, and robustness against shortcut behaviors. We apply strict filtering to ensure trajectory quality, removing trivial cases and enforcing a consistent reasoning and tool-use format. To avoid shortcut learning such as lucky guesses based on partial evidence, we require trajectories contains explicitly verification of all conditions specified in the query. We further retain trajectories that involve long and iterative exploration by compressing earlier steps, allowing long-horizon reasoning to be preserved without length constraints. Finally, we scale data collection and increase behavioral diversity by reusing rollout trajectories from subsequent reinforcement learning stages.
@@ 3.2 Scalable Asynchronous Agentic RL Framework
Compared with single-turn scenarios such as standard reasoning tasks, agentic training involves multi-turn interactions with variable environments (ENV) or tools, which pose new challenges to RL Infrastructure. During a typical multi-turn rollout stage, it repeatedly interleaves the LLM generation, environment execution, and reward assessment to construct the final trajectory for RL training. In this setting, trajectories are not only long-tailed skewed but also involve unpredictable and latency-skewed environment interactions, which lead to device underutilization under batch setting (Fu et al., 2025; Zhang et al., 2025a; Lu et al., 2025b). Additionally, our production cluster consists of mid-range accelerators, especially with only around 60GB of available device memory. The hardware constraints, together with its software ecosystem, pose significant challenges to achieving stable and scalable agentic RL training. To address these problems, we extend our multi-version asynchronous training system, DORA (Dynamic ORchestration for Asynchronous rollout) (Team et al., 2025), to fully support large-scale RL training in multi-turn agentic scenarios.
@@
To minimize device idleness in the agentic setting, we introduce a fully streaming asynchronous pipeline both within the rollout process and between rollout and training, based on streaming RPC (Team et al., 2025). Within the rollout loop in the RolloutManager, we remove batch barriers, enabling LLM generation, environment execution, and reward computation to be executed on remote workers at the granularity of individual samples. This prevents accelerators from idling while waiting for batched ENV calls to complete. To further address the long-tailed generation problem with training stability, our DORA system enables multi-version asynchronous training, where trajectories generated by different model versions are immediately enqueued upon completion. Within a step, our multi-version generation instances continue rollout using multiple previous model versions, while the Trainer can initiate training as soon as its conditions are met, or elastically scale up additional generation instances for free extra throughput when training devices are idle.
@@ 7 Conclusion
Dengchang Zhao Jiarui Zhao Shijun Dai Shizhe Wu Yang Bai
Yuxin Chen Di Xiu Jingang Wang Shuaikang Liu Yang Liu
Yuxin Liu Dishan Liu Jinluan Yang Shuang Zhou Yang Yang
Zeyang Yu Dongyu Ru Jinrui Ding Shuo Wang Yang Zheng
Zhao Yang Dunwei Tu Jinwei Xiao Songyuan Zhao Yanyu Chen
Zhengsheng Huang Fan Wu Jiyuan He Tao Liang Yaoming Wang
Zhengyu Chen Fengcheng Yuan Juncan Xu Tianhao Hu Yaoming Zhu
Zhijian Liu Fengcun Li Kefeng Zhang Tianze Chen Yaorui Shi
Zhikang Xia Gang Xu Keheng Wang Wei Liu Yaqi Huo
Zhimin Lin Guanyu Wu Li Wei Wei Shi Yerui Sun
Zhiyuan Yao Guoyuan Lin Lianhui Ma Wei Wang Yi Zhang
Zhuofan Chen Haibin Wang Lin Qiu Weifeng Tang Yifan Lu
Zhuowen Han Hansi Yang Lingbing Kong Wenjie Shi Yifan Zhao
Zijian Zhang Hao Yang Lingchuan Liu Wenlong Zhu Yihao Chen
Ziran Li Haonan Yan Linsen Guo Wentao Chen Yi-Kai Zhang
Ziwen Wang Haoxiang Ma Mengshen Zhu Wentao Shi* Yitao Zhai
Ziyuan Zhuang Haoxing Wen Mengxia Shen Xi Su Yongjing Yin
LongCat-Flash